Circular No. 6046

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN SCORPIUS S. N. Zhang, C. A. Wilson, B. A. Harmon, G. J. Fishman, R. B.

Wilson, W. S. Paciesas, M. Scott, and B. C. Rubin report for the Compton Gamma Ray Observatory BATSE team: “An x-ray nova has been observed at α = 16^{h}55^{m}, δ = −40^{o}.5 (equinox 2000.0; error radius 0^{o}.3). The source (GRO J1655–40) first appeared in data from July 27 and has increased to its present level of ∼ 1.1 Crab (20–100 keV) on Aug. 1. Significant flux is observed to 200 keV. A power law with a spectral index of −3.15 ± 0.07 fits the data well from 20 to 200 keV. Pulsed emission is not seen in the period range from 3 to 300 s at a limiting sensitivity of ∼ 10 percent of the steady flux. No previously-cataloged x-ray source is known at this location. It has also been determined that the nearby x-ray pulsar OAO 1657–415, currently observed by BATSE, is not responsible for the observed emission.”

SUPERNOVA 1994W IN NGC 4041 A. V. Filippenko and A. J. Barth, University of California at Berkeley,

report: “A CCD spectrogram (range 425-704 nm, resolution 0.7 nm), obtained on Aug. 4 with the Shane 3-m reflector at Lick Observatory, confirms that SN 1994W is a peculiar type-II supernova (cf. IAUC 6044). Relatively narrow (FWHM = 1200 km/s) hydrogen Balmer emission lines are superposed on much broader bases (FWZI = 7500 km/s); broad (FWHM ∼ 5000 km/s) He I 587.6-nm emission is also visible, with no narrow component. Narrow Fe II emission lines are weaker but present. Unlike the case in normal type-II supernovae, the Balmer lines do not show broad P-Cyg absorption. However, they do exhibit narrow (FWHM = 300 km/s) P-Cyg absorption components with minima displaced by 700 km/s from the emission-line cores. Similar features are seen in some of the Fe II lines, especially those of multiplet 42; they may be produced by a wind from the progenitor. Previous supernovae showing such characteristics have often been strong radio sources, due to interaction of the ejecta with dense circumstellar gas. Further observations at all wavelengths are encouraged. Note that the offsets reported by Richmond et al. (cf. IAUC 6043) were based on a slightly incorrect value for the pixel scale; SN 1994W is actually 7′′.3 west and 17′′.7 north of the nucleus, consistent with the values given by Pollas (ibid.).”

1994 August 4 Daniel W. E. Green